OpenStudy (anonymous):

I need help, I'll fan and medal Which literal equation is equivalent to 1.4 = kx A. k= x divided by 1.4 B. x= 1.4 multiplied by k C. k= 1.4 divided by x D. x= k divided by 1.4 Please explain how to get the answer I really don't get it.

OpenStudy (alexandervonhumboldt2):

lets divide both sides by x: so we will get 1.4 = kx => kx/x=1.4/x => k=1.4/x

OpenStudy (alexandervonhumboldt2):

lets divide both sides by k: kx/k=1.4/k => x=1.4/k

OpenStudy (anonymous):

Why do I divide k and x on both sides? And how is that possible

OpenStudy (alexandervonhumboldt2):

well so in basic form: ab=c => a=c/b b=c/a ab-c=0